### v1

# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our strong negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further study the impact of different negative sample construction methods and the number of negative samples on the geometric reasoning performance of LMM, yielding fruitful conclusions. The code and dataset are available at https://github.com/THU-KEG/MMGeoLM.
Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
Kai Sun, Yushi Bai, Zhen Yang, Jiajie Zhang, Ji Qi, Lei Hou, Juanzi Li Tsinghua University
11footnotetext: Equal contribution
1 Introduction
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (LMMs) Anthropic (2024); OpenAI (2023); Bai et al. (2023). It requires fine-grained identification of visual elements Lu et al. (2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them He et al. (2024).
[…]
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP Patel et al. (2024); Yang et al. (2023); Goel et al. (2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o OpenAI (2024a), Claude-3 Anthropic (2024), and Qwen2.5-VL Bai et al. (2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
To address these shortcomings, recent research has focused strategies such as fine-tuning on specialized mathematical datasets Gao et al. (2023); Zhang et al. (2024c); Peng et al. (2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs Qi et al. (2024); Wang et al. (2024) by aligning images with corresponding captions. However, the construction of image-caption pairs is inherently susceptible to hallucinations, as many captions are generated by existing LMMs. As a result, training with only these positive samples may lead to spurious alignment Zhang et al. (2024b). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Model Training Training Data MM-MATH
AltCLIP No No 23.8
AltCLIP In-Batch Randomly-sampled Negative (400K) 24.6
AltCLIP MMCLIP ①: Retrieval-based Negative (100K) 26.6
AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
AltCLIP MMCLIP ③: Image Negative (4K) 28.2
AltCLIP MMCLIP ①+② 28.4
AltCLIP MMCLIP ①+②+③ 30.1
Table 1: Comparison of geometric reasoning performance on MM-MATH across LMMs with vision encoders trained with different hard negative data.
In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in the vision encoder. For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives. For image-based contrastive learning, we introduce a novel method that leverages a large language model (LLM) to generate a detailed caption and corresponding diagram generation code for a given geometry problem, forming the positive image sample. The LLM then perturbs the code to create visually similar but geometrically incorrect diagrams, which serve as hard negative samples. Additionally, we modify the original CLIP training framework and propose MMCLIP, a method designed to handle an arbitrary number of hard negative samples centered around a single image or caption. Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
We conduct extensive experiments to evaluate the effectiveness of our CLIP training for LMMs. Specifically, we assess our trained MMGeoLM across four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVISTA, and WE-MATH) and open-ended questions (MM-MATH). Our results show that the proposed model outperforms all existing open-source models on GeoQA and MathVISTA, and achieves state-of-the-art performance on MM-MATH, surpassing GPT-4o by 7.5%. We further analyze the impact of different types of hard negative samples. Notably, we find that negative samples constructed from authentic, exam-based images significantly boost model performance—using just 4K image negatives yields better results than over 100K text negatives. Lastly, we observe that increasing the number of hard negatives improves performance up to a certain threshold, beyond which it leads to diminishing returns and eventual degradation.
Figure 2: Image-based and text-based hard negative construction and the corresponding MMCLIP training method.
2 Method
In this section, we first introduce our proposed geometric diagram strong negative construction method in 2.1. Subsequently, we present the corresponding negative captions approach for geometric elements in 2.2. Finally, we describe a training strategy of handling an arbitrary number of negative examples in 2.3, followed by our newly collected multimodal mathematical dataset in 2.4.
2.1 Negative Images Construction
We construct negative geometric examples by generating geometric diagram codes for positive images and perturbing them to create negatives. Existing generation methods Zhang et al. (2024c); Zou et al. (2024); Wei et al. (2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic exam questions, created by educators using professional geometry software McClintock et al. (2002), produce diagrams that are more diverse and accurate. Specifically, in these authentic geometry problems Sun et al. (2024), geometric elements are often described explicitly (e.g., AB is 8 cm) or implicitly through reasoning (e.g., using the Pythagorean theorem). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs OpenAI (2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional informative annotations (see Appendix A for prompt details).
To handle occasional errors in model-generated code, we adopt a lightweight model GLM et al. (2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro DeepMind (2025), avoiding noise from question-based captions. Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, We design prompts (Appendix A) for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
2.2 Negative Captions Construction
We propose two methods for constructing text-based hard negative samples. The first method retrieves captions similar to each positive example from a large image-caption dataset. The second method generates targeted hard negative samples based on reasoning errors observed during LMM inference. Some examples of these hard negative samples are shown in Figure 1.
Retrieval-based Hard Negative.
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance Karpukhin et al. (2020); Xiong et al. (2020); Huang et al. (2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives Xiong et al. (2020); Zhou et al. (2022); Yang et al. (2024). To address this issue, we build upon the recently proposed MAVIS dataset Zhang et al. (2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS Zhou et al. (2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
Rule-based Hard Negative.
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships Yuksekgonul et al. (2022); Zhang et al. (2024a), while others leverage in-context learning to generate meaningful modifications Patel et al. (2025). However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level without considering downstream reasoning errors. For example, as shown in Figure 1, LMMs incorrectly interpret the parallel line $AB$ and $CD$ as forming triangle $ABC$, leading to errors in solving the geometric problem. Simply altering positive captions without considering LMM reasoning errors provides limited benefits in improving geometric element recognition.
To address this issue, we analyze the evaluation results of LMMs on MM-MATH Sun et al. (2024) and identify four major types of image element recognition errors in their responses. Based on these error types, we design corresponding rule-based strategies to construct hard negative samples:
  1. 1.
    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ are invalid).
  2. 2.
    Shape attributes: Altering properties such as changing squares to rectangles or right triangles to isosceles triangles.
  3. 3.
    Geometric relationships: Modifying relationships such as parallelism between two lines or similarity between two triangles.
  4. 4.
    Numerical values: Adjusting numerical values in captions, such as modifying angles or segment lengths.
We use GLM-4 GLM et al. (2024) to generate 10 rule-based hard negatives for each positive example. An example is shown in Figure 2.
2.3 Hard Negative CLIP Training
Though CLIP-trained vision encoders are widely used in LMMs Radford et al. (2021); Doveh et al. (2023b), they typically adopt in-batch negative sampling during training, which limits their capacity for fine-grained image understanding. While methods like NegCLIP Yuksekgonul et al. (2022) introduce hard negatives, they still rely heavily on in-batch samples. Considering the methods in Section 2.2 have allowed us to construct a large number of hard negatives for each image, we propose MMCLIP training strategy, which focuses on each single image and only uses its corresponding hard negatives for training, entirely abandoning traditional in-batch negatives. In this way, the vision encoder is forced to distinguish subtle differences in all cases, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive example and $T_{j}^{-}$ is the negative samples conducted around the image, using the methods in Section 2.2. Our loss function is formulated as
$\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings of the input instances. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ are expected to be higher than traditional in-batch negatives. From the above equation, we calculate the gradient for positive and negative samples:
$\frac{\partial\mathcal{L}}{\partial s_{i}^{+}}=-\frac{\sum_{j=1}^{N}\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (2)
$\frac{\partial\mathcal{L}}{\partial s^{-}_{j}}=\frac{\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (3)
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. We further validate the effectiveness of hard negatives and analyze the impact of their quantity in our experiments in Section 4.2.
2.4 Supervised Fine-tuning Data
[…]
3 Multimodal Mathematical Training
3.1 Architecture
We adopt the LLaVA architecture Liu et al. (2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use Mammoth2-7B Yue et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025). The vision encoder is based on AltCLIP Chen et al. (2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
3.2 Training Details
In the vision-text alignment stage, we first pre-train AltCLIP on the 400K MAVIS image-caption alignment dataset Zhang et al. (2024c) to improve the model’s generalization for geometric problems. Next, we fine-tune the model following the MMCLIP training strategy, using 100K hard caption negative samples, each consisting of 10 rule-based and 30 retrieval-based negatives to enhance the model’s understanding of geometric elements. In addition, we fine-tune the model on 4K hard image negative samples, where each example includes 10 negative geometric diagrams. In the second stage, we tune the MLP adapter using 67K G-LLAVA Gao et al. (2023) image-text alignment data. In the third stage, we perform supervised fine-tuning on a combined dataset comprising 300K MAVIS instruction data Zhang et al. (2024c), 117K G-LLaVA instruction data Gao et al. (2023), and 17K open-ended geometric problems. The open-ended dataset includes 12K problems from our collected dataset and 5K randomly sampled MM-MATH Sun et al. (2024) geometric problems, with the rest 700 problems reserved for evaluation. We name the trained model as MMGeoLM.
4 Experiments
To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs. Additionally, we analyze the impact of different vision encoder types and training strategies, as well as the number of hard negatives in MMGeoLM.
4.1 LMM Geometric Reasoning Evaluation
4.1.1 Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA Chen et al. (2021), a geometric question-answering task based on plane geometry; We-Math Qiao et al. (2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista Lu et al. (2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-MATH Sun et al. (2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-MATH, we evaluate models across easy, medium, and hard categories using a test set of 700 problems excluded from training. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
4.1.2 Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus Anthropic (2024), Claude-3.5-Sonnet, GPT-4o OpenAI (2024a), and GPT-4V. Open-source LMMs: MAVIS-7B Zhang et al. (2024c), G-LLaVA-7B Gao et al. (2023), InternVL2-8B Chen et al. (2024), LLaVA-OneVision-7B Li et al. (2024), Qwen2.5-VL-7B-Instruct Bai et al. (2025), Chimera-Reasoner-8B Peng et al. (2024b), InternLM-XComposer2-7B Dong et al. (2024), Phi-3-Vision-128K-Instruct Abdin et al. (2024), and Math-LLaVA-13B Shi et al. (2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks Chen et al. (2021); Lu et al. (2023) since 2023. Additionally, we include human evaluation baselines, using scores from prior studies Zhang et al. (2024c); Sun et al. (2024).
Model GeoQA MathVista We-Math MM-MATH
GPS ALG S2 S3 Easy Med. Hard Avg
Human* 92.3 48.4 50.9 - - 90.7 81.9 47.6 80.4
Claude-3-Opus* 44.5 46.3 46.6 32.9 23.0 29.5 19.3 3.6 20.3
Claude-3.5-Sonnet 65.1 66.3 68.4 64.7 62.1 34.4 31.9 13.6 31.7
GPT-4V* - 50.5 53.0 49.2 38.2 37.8 21.2 1.8 23.1
GPT-4o 58.9 62.7 65.4 58.0 43.6 45.8 30.0 10.9 31.8
MAVIS-7B* 68.3 64.1 59.2 37.9 34.6 - - - -
G-LLaVA-7B* 67.0 56.7 - 30.1 32.7 - - - -
Math-LLaVA-13B 48.4 55.6 55.0 31.7 23.0 - - - -
InternVL2-8B 56.4 60.5 60.8 41.1 37.1 33.6 21.7 9.0 23.3
LLaVA-OneVision-7B 64.6 54.6 53.1 28.7 22.3 40.5 24.8 10.5 27.0
Chimera-Reasoner-8B 69.6 48.5 42.6 29.9 31.6 36.2 22.2 9.0 24.1
Phi-3-Vision-128K-Instruct 29.5 38.8 40.0 33.3 30.1 12.9 7.1 0.0 7.8
Qwen2.5-VL-7B-Instruct 59.0 67.7 66.9 58.1 50.6 50.7 32.2 14.2 34.8
InternLM-XComposer2-7B 38.8 63.0 56.6 33.1 33.0 18.9 12.2 4.5 13.1
MMGeoLM-Mammoth2-7B 68.5 68.7 69.5 41.5 37.5 52.5 32.4 4.5 36.9
   w/o MMCLIP 55.4 52.2 51.3 34.1 35.3 42.5 21.8 4.5 26.8
   Original AltCLIP 46.7 45.8 46.8 30.1 29.2 38.2 20.4 1.6 25.9
MMGeoLM-Qwen2.5-7B 69.2 69.8 68.1 39.8 36.4 55.3 36.9 9.0 39.2
   w/o MMCLIP 56.3 54.1 50.3 35.2 35.1 50.5 28.8 4.5 34.3
   Original AltCLIP 50.7 47.6 48.9 33.1 32.7 52.5 30.8 4.5 33.7
Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and MM-MATH. ‘S2/S3’ denote the two-step settings in We-Math. Results marked * are taken from their original papers.
4.1.3 Overall Results
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-MATH benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA Peng et al. (2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-Mammoth2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-Mammoth2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-MATH benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the original AltCLIP and without MMCLIP (AltCLIP trained only with in-batch learning on MAVIS), the vision encoder trained with our constructed hard negative samples and the MMCLIP training method significantly enhances MMGeoLM’s geometric understanding capabilities.
Figure 3: Performance evaluation of MMGeoLMs built with six AltCLIP trained on different contrastive datasets.
Figure 4: Comparison of two image negative strategies: Mavis uses 100K samples from the Mavis dataset, while MM-MATH-Align uses our 4K constructed samples.
Figure 5: Retrieval performance of AltCLIP variants trained with different strategies. ‘Random + X’ means sequential training with random negatives followed by negatives created with the X strategy.
4.2 Experiments on Vision Encoder
We further quantitatively evaluate the performance of various vision encoders trained with different datasets and strategies. The experiments are divided into three parts: (1) performance evaluation of six LMMs trained with different AltCLIP variants across two benchmarks, (2) top-1 positive example retrieval performance evaluation of six AltCLIP models, and (3) varying number of hard caption negatives.
4.2.1 Setup
We compare the following six types of AltCLIP models are as follows:
  • Original AltCLIP: No additional data training.
  • Random AltCLIP: Trained with randomly selected negative samples within a batch (batch size = 10) using 400K MAVIS data.
  • Retrieved AltCLIP: Trained with 100K samples, each with 10 retrieval-based caption negatives.
  • Rule-based AltCLIP: Trained with 100K samples, each with 10 rule-based caption negatives.
  • Negative-images AltCLIP: Trained with 4K samples, each with 10 hard image negatives.
  • Retrieved + Rule-based + Negative-images: Trained on all three datasets combined.
The Random AltCLIP follows OpenAI’s training approach Radford et al. (2021), while the Retrieved, Rule-based, Negatives-images and Retrieved + Rule-based + Negatives-images AltCLIP models employ our proposed MMCLIP training approach. In the second and third stages, LMMs are trained on 67K G-LLaVA alignment data and 17K open-ended geometric reasoning problems.
For the retrieval experiments, we assess the retrieval performance of trained AltCLIP models by selecting positive samples from a large pool of negative samples. Retrieval performance is evaluated using the Hit@1 metric. We construct four types of negative datasets: Random Negative: Caption negatives randomly selected in a training batch. Retrieval Negative: Caption negatives retrieved using the SimANS Zhou et al. (2022) model, selecting the top 100 most similar captions to the positive sample. Rule-based Negative: Caption negatives generated according to the rules defined in section 2.2. Image Negative: Image negatives generated from real-world geometric problems by MM-MATH.
For the number of negative experiments, we investigate the impact of hard caption negative samples on MMGeoLM varying from 5 to 50.
4.2.2 Results
As shown in Figure 4, LMMs trained with hard negative samples consistently improve performance on the GeoQA and MM-MATH benchmarks. Notably, on MM-MATH, Negative-images AltCLIP, trained with only 4K image negative samples, surpasses Retrieved AltCLIP and Rule-based AltCLIP by 0.2% and 1.7%, respectively. On MM-MATH, Rule-based, Retrieved, and Image-negative AltCLIP consistently outperform Random AltCLIP on easy and medium-level questions. These results confirm the effectiveness of both hard negative construction methods. Furthermore, AltCLIP trained with a combination of caption and image negatives achieves the best performance on both GeoQA and MM-MATH, demonstrating their complementary effect in improving model performance. Further experiments on image negatives (Figure 4) reveal that Negative-image-MM-MATH-Align, constructed from real-world images, outperforms Negative-image-Mavis. This suggests that high-quality, realistic image negatives have a more pronounced impact on model performance than large-scale synthetic counterparts.
The results of the retrieval performance experiments are shown in Figure 5. Under the negative-images setting, Negative-images AltCLIP achieves a retrieval score of 23%, while other AltCLIP variants remain below 15%. These results highlight the difficulty of retrieving realistic images. Notably, the dataset includes only 4,021 image negative samples, which is likely insufficient to fully exploit the model’s performance potential. In the retrieval-based and rule-based negative settings, Random AltCLIP outperforms Rule-based AltCLIP in retrieval-based settings and exceeds Retrieved AltCLIP in rule-based settings. These results indicate that while hard negative training enhances performance within specific domains, it may limit generalization. Furthermore, pre-training on random negatives followed by hard negatives improves both generalization and domain-specific performance, ultimately enhancing overall model effectiveness.
Figure 6: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative samples.
As shown in Figure 6, MMGeoLM’s performance on MM-MATH improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point. These results indicate that hard caption negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance. Training with only five hard negative samples outperforms training with ten randomly selected negatives (as shown in Table 1), demonstrating that hard negatives enhance multimodal mathematical reasoning more effectively.
5 Related Work
Mathematics-related research has recently received significant attention in LLMs and LMMs. For text-only mathematical reasoning, several works employ external tools like Tora Gou et al. (2023) or intermediate step decomposition methods such as MAmmoTH Yue et al. (2023), Metamath Yu et al. (2023) and Math-Shepherd Yue et al. (2023). For multimodal mathematics tasks involving images, benchmarks have been proposed, including mathematical competition-oriented OlympiadBench He et al. (2024), geometry-focused datasets like Geometry3K Lu et al. (2021), VisScience Jiang et al. (2024), UniGeo Chen et al. (2022a) and GeoQA Chen et al. (2022a). Many works take effort on training multimodel math, including G-llava Gao et al. (2023), Meta-LLaVA Shi et al. (2024) and MAVIS Zhang et al. (2024c). However, these multimodal approaches typically suffer from inaccuracies in recognizing geometric elements, where our paper specifically addresses the enhancement of visual recognition accuracy.
Vision encoders play a crucial role in understanding image content. While previous works Zhang et al. (2024a); Doveh et al. (2023a, b); Singh et al. (2023) have leveraged negative captions to enhance image comprehension, but not address the recognition of geometric elements within images. Methods such as NegCLIP Yuksekgonul et al. (2022) and TriCLIP Yang et al. (2024) focus on learning from negative samples, but their applicability is limited by the constrained number of negative samples. In this work, we propose a robust negative learning strategy that explicitly constructs negative samples for geometric elements while enabling arbitrary expansion, thereby improving geometric understanding in multimodal tasks.
6 Conclusion
This paper introduced MMCLIP, a novel hard negative contrastive learning framework to enhance LMMs’ geometric reasoning by refining their vision encoders with systematically generated image and text-based negatives. Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks. This work demonstrates the power of tailored hard negative sampling for developing LMMs with fine-grained visual understanding in specialized domains such as geometric understanding and reasoning.
Limitations
Despite the effectiveness of the proposed image hard negatives constructed by perturbing Python scripts, the method heavily relies on the accuracy of LLM-generated code and captions. Any systematic biases in these synthetic constructions may introduce artifacts that differ from human-designed geometric problems. As a result, the model’s performance on real-world visual inputs remains uncertain and requires further validation on diverse real-world datasets.
Appendix A Prompts
After multiple experimental iterations, we constructed two different types of prompts for our experiments, as illustrated in Figures 7 and 8. Specifically, Figure 7 demonstrates the prompt constructed by inputting the textual description and answer of realistic geometry problems to obtain the corresponding geometric diagram Python script. Meanwhile, Figure 8 illustrates the prompt obtained by modifying captions from positive examples to create slightly perturbed negative samples.
Figure 7: Python script prompt from question and true answer in geometric problems
Figure 8: Negative caption generation prompt from positive captions
Appendix B Figure Comparison
Figure 9 presents illustrative examples of image negative sample, covering analytical geometry and planar geometry. Geometric elements include triangles, quadrilaterals, as well as various relational properties such as perpendicularity and intersection. Additionally, certain metric properties absent from the original diagrams are explicitly indicated. Specifically, the first column shows the original diagrams, and the second column depicts diagrams generated from the given questions and the true answer, which, despite slight discrepancies, accurately capture the overall outlines of the original diagrams. The third and fourth columns display the constructed negative diagram instances—generated based upon their corresponding positive examples—demonstrating notable similarities but also clear and meaningful distinctions.
Figure 9: Comparison of generated geometric figures. The first column shows the original images, the second column presents the positive samples, and the last two columns illustrate the constructed negative samples.
Appendix C Case Study
Figure 10: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. MMGeoLM correctly answers the problem using a different solution compared to True Answer, while GPT-4o generates a non-existent element $\triangle ABD$.
Figure 11: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
C.1 Case Study
In Figures 10 and 11, we compare the solutions generated by GPT-4o and our MMGeoLM. In Figure 10, the red text highlights GPT-4o’s misrecognition of two triangles in the image, leading to an incorrect final result. In contrast, MMGeoLM produces a correct solution, albeit through a reasoning process different from the ground-truth answer. This highlights MMGeoLM’s capability to generate diverse problem-solving approaches. Figure 11 illustrates errors in image-based reasoning for both MMGeoLM and GPT-4o. While both models make mistakes, MMGeoLM’s reasoning aligns more closely with the correct solution, demonstrating improved mathematical reasoning. Overall, our trained MMGeoLM model demonstrates greater reasoning diversity and enhanced geometric problem-solving performance compared to current mainstream models.