### v2

# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our hard negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations. These analyses yield important insights into optimizing hard negative strategies for geometric reasoning tasks. https://github.com/THU-KEG/MMGeoLM.
Introduction
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic 2024; OpenAI 2023; Bai et al. 2023). It requires fine-grained identification of visual elements (Lu et al. 2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al. 2024).
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al. 2024; Yang et al. 2023; Goel et al. 2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI 2024a), Claude-3 (Anthropic 2024), and Qwen2.5-VL (Bai et al. 2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
[…]
To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al. 2023; Zhang et al. 2024b; Peng et al. 2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al. 2024; Wang et al. 2024) by aligning images with corresponding captions. However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al. 2024a; Doveh et al. 2023a). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Model Training Training Data $M$-Math
AltCLIP No No 23.8
AltCLIP In-Batch Randomly-sampled Negative (400K) 24.6
AltCLIP MMCLIP ①: Retrieval-based Negative (100K) 26.6
AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
AltCLIP MMCLIP ③: Image-based Negative (4K) 28.2
AltCLIP MMCLIP ①+② 28.4
AltCLIP MMCLIP ①+②+③ 30.1
Table 1: Comparison of geometric reasoning performance on MM-Math across LMMs with vision encoders trained with different hard negative data.
In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in vision encoder. For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives. For image-based contrastive learning, we introduce a novel method that leverages a large language model (LLM) to generate a detailed caption and corresponding diagram generation code for a given geometry problem, forming the positive image sample. The LLM then perturbs the code to create visually similar but geometrically incorrect diagrams, which serve as hard negative samples. Additionally, we modify the original CLIP training framework and propose MMCLIP, a method designed to handle an arbitrary number of hard negative samples centered around a single image or caption. Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
We evaluate our trained MMGeoLM on four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVista, and We-Math) and open-ended questions (MM-Math). Results show that the proposed model outperforms all existing open-source models on GeoQA and MathVista, and achieves state-of-the-art performance on MM-Math, surpassing GPT-4o by 7.4%. Ablation studies confirm both text-based and image-based hard negatives benefit model performance. Notably, using only 4K image-based negatives yields better results than 100K retrieval-based ones. These findings demonstrate that our hard negative construction and training strategy significantly enhances reasoning accuracy in geometric reasoning.
Figure 2: Image-based and text-based hard negative construction and the corresponding MMCLIP training method.
Method
Negative Images Construction
We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives. Existing generation methods (Zhang et al. 2024b; Zou et al. 2024; Wei et al. 2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic geometry problems (Sun et al. 2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock, Jiang, and July 2002). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al. 2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind 2025). Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
Negative Captions Construction
We propose two methods for constructing text-based hard negative samples. The first method retrieves captions similar to each positive example from a large image-caption dataset. The second method generates targeted hard negative samples based on reasoning errors observed during LMMs’ inference.
Retrieval-based Hard Negative.
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al. 2020; Xiong et al. 2020; Huang et al. 2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al. 2020; Zhou et al. 2022; Yang et al. 2024). To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al. 2024b), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS (Zhou et al. 2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
Rule-based Hard Negative.
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al. 2022; Zhang, Awal, and Agrawal 2024), while others leverage in-context learning to generate meaningful modifications (Patel et al. 2025) . However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al. 2024) and identify four major types of image element recognition errors:
  1. 1.
    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ is invalid).
  2. 2.
    Shape attributes: Altering properties such as changing squares to rectangles or right triangles to isosceles triangles.
  3. 3.
    Geometric relationships: Modifying relationships such as parallelism between two lines or similarity between two triangles.
  4. 4.
    Numerical values: Adjusting numerical values in captions, such as modifying angles or segment lengths.
To address these error types, we design rule-based strategies that use GLM-4 (GLM et al. 2024) to modify each positive caption and generate 10 hard negative examples.
These text-based negatives are used in model training, and two representative types are shown in Figure 2.
Hard Negative CLIP Training
Figure 3: Overview of the MMGeoLM training pipeline, including main and each ablation experiment configurations, training strategies, and datasets.
Though CLIP-trained vision encoders are widely used in LMMs (Radford et al. 2021; Doveh et al. 2023b), they typically combine random examples for a batch, which limits their capacity for fine-grained image understanding during training. While methods like NegCLIP (Yuksekgonul et al. 2022) introduce hard negatives, they still rely heavily on randomly selected negatives within a batch. Considering the methods in the above section have allowed us to construct a large number of hard negatives, we propose a MMCLIP training strategy, which focuses on each single sample and only uses its corresponding hard negatives for training. In this way, the vision encoder is forced to distinguish subtle differences, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a text-based negative training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive samples and $T_{j}^{-}$ is the negative samples conducted around the image. Our loss function is formulated as follows:
$\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ is expected to be higher than a randomly chosen one. From the above equation, we calculate the gradient for positive and negative samples:
$\frac{\partial\mathcal{L}}{\partial s_{i}^{+}}=-\frac{\sum_{j=1}^{N}\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (2)
$\frac{\partial\mathcal{L}}{\partial s^{-}_{j}}=\frac{\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (3)
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. Unlike in-batch training, which relies on batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for rule-based or retrieval-based unimodal negatives. We further validate different training strategies only for hard negatives in the ablation experiments.
Supervised Fine-tuning Data
[…]
Multimodal Mathematical Training
Architecture
We adopt the LLaVA architecture (Liu et al. 2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use MAmmoTH2-7B (Yue et al. 2024) and Qwen2.5-7B-Instruct (Qwen et al. 2025). The vision encoder is based on AltCLIP (Chen et al. 2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
Training Details
Model GeoQA MathVista We-Math $M$-Math
GPS ALG S2 S3 Easy Med Hard Avg
Human* 92.3 48.4 50.9 - - 90.7 81.9 47.6 80.4
Claude-3-Opus* 44.5 46.3 46.6 32.9 23.0 29.5 19.3 3.6 20.3
Claude-3.5-Sonnet 65.1 66.3 68.4 64.7 62.1 34.4 31.9 13.6 31.7
GPT-4V* - 50.5 53.0 49.2 38.2 37.8 21.2 1.8 23.1
GPT-4o 58.9 62.7 65.4 58.0 43.6 45.8 30.0 10.9 31.8
MAVIS-7B* 68.3 64.1 59.2 37.9 34.6 - - - -
G-LLaVA-7B* 67.0 56.7 - 30.1 32.7 - - - -
Math-LLaVA-13B 48.4 55.6 55.0 31.7 23.0 - - - -
InternVL2-8B 56.4 60.5 60.8 41.1 37.1 33.6 21.7 9.0 23.3
LLaVA-OneVision-7B 64.6 54.6 53.1 28.7 22.3 40.5 24.8 10.5 27.0
Chimera-Reasoner-8B 69.6 48.5 42.6 29.9 31.6 36.2 22.2 9.0 24.1
Phi-3-Vision-128K-Instruct 29.5 38.8 40.0 33.3 30.1 12.9 7.1 0.0 7.8
Qwen2.5-VL-7B-Instruct 59.0 67.7 66.9 58.1 50.6 50.7 32.2 14.2 34.8
InternLM-XComposer2-7B 38.8 63.0 56.6 33.1 33.0 18.9 12.2 4.5 13.1
MMGeoLM-MAmmoTH2-7B 68.5 68.7 69.5 41.5 37.5 52.5 32.4 4.5 36.9
  w/o MMCLIP 55.4 52.2 51.3 34.1 35.3 42.5 21.8 4.5 26.8
  Original AltCLIP 46.7 45.8 46.8 30.1 29.2 38.2 20.4 1.6 25.9
MMGeoLM-Qwen2.5-7B 69.2 69.8 68.1 39.8 36.4 55.3 36.9 9.0 39.2
  w/o MMCLIP 56.3 54.1 50.3 35.2 35.1 50.5 28.8 4.5 34.3
  Original AltCLIP 50.7 47.6 48.9 33.1 32.7 52.5 30.8 4.5 33.7
Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and $M$-Math. ‘S2/S3’ denotes the two-step settings in We-Math. Results marked * are taken from their original papers.
Our training pipeline, illustrated in Figure 3, consists of three stages covering both main and ablation experiments. In the first stage, we perform vision-text alignment training. For text-based hard negatives, we use the Mavis image-caption alignment dataset (Zhang et al. 2024b). For image-based negatives, we collect 4K geometry questions from real middle-school exams. Additional details, including the types of negatives and the negative-to-positive ratio, are provided in Figure 3. In the second stage, we tune the MLP adapter using 67K image-text pairs from the G-LLaVA dataset (Gao et al. 2023). In the third stage, we conduct supervised fine-tuning. For main experiments, we use a combined dataset comprising 300K Mavis instruction data (Zhang et al. 2024b), 117K G-LLaVA instruction data (Gao et al. 2023), and 17K open-ended geometry questions. For ablation studies, we only use the 17K newly collected geometry problems for efficient validation. We name the final trained model as MMGeoLM.
Experiments
To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs. Additionally, we analyze the impact of different vision encoder types, image-based negatives and training strategies in the subsequent ablation studies.
LMM Geometric Reasoning Evaluation
Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA (Chen et al. 2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al. 2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al. 2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-Math (Sun et al. 2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic 2024), Claude-3.5-Sonnet, GPT-4o (OpenAI 2024a), and GPT-4V. Open-source LMMs: MAVIS-7B (Zhang et al. 2024b), G-LLaVA-7B (Gao et al. 2023), InternVL2-8B (Chen et al. 2024), LLaVA-OneVision-7B (Li et al. 2024), Qwen2.5-VL-7B-Instruct (Bai et al. 2025), Chimera-Reasoner-8B (Peng et al. 2024b), InternLM-XComposer2-7B (Dong et al. 2024), Phi-3-Vision-128K-Instruct (Abdin et al. 2024), and Math-LLaVA-13B (Shi et al. 2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al. 2021; Lu et al. 2023). Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al. 2024b; Sun et al. 2024).
Overall Results
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA (Peng et al. 2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-MAmmoTH2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-MAmmoTH2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-Math benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the w/o MMCLIP (AltCLIP trained only with random negatives in a batch), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
Ablation Study 1: Different Negatives Types
Vision Encoder #Num. GeoQA $M$-MATH
Easy Med Hard Avg
Original - 45.4 40.5 20.8 4.5 23.8
$\text{Random}_{\text{10}}$ 400K 45.0 39.8 22.7 4.5 24.9
$\text{Retrieval}_{\text{10}}$ 100K 53.0 43.6 24.8 4.5 26.6
$\text{Rule}_{\text{10}}$ 100K 48.8 45.4 27.6 4.5 28.1
$\text{Image-based}_{\text{10}}$ 4K 54.9 45.6 26.5 4.5 29.0
$\text{Retrieval}_{\text{10}}$$\text{Rule}_{\text{10}}$ 200K 56.5 47.5 28.3 4.5 29.4
All Negatives 204K 58.2 49.8 29.1 4.5 30.4
Table 3: Ablation study on vision encoder training with different types and quantities of hard negatives. The bottom panel reports performance when combining multiple types.
To further evaluate the performance of vision encoders trained with different types of hard negatives, we compare seven negatives summarized in Table 3. All models are based on AltCLIP. The Original vision encoder is directly initialized without any additional training, while $\text{Random}_{\text{10}}$ follows the conventional in-batch random negatives sampling strategy (Radford et al. 2021). The remaining five vision encoders are trained with our proposed MMCLIP method. The subscript 10 indicates a 1:10 positive-to-negative ratio.
Results
As shown in Table 3, vision encoders trained with hard negatives consistently outperform random in-batch negatives. even with only 4K image-based negatives , the $\text{Image-based}_{\text{10}}$ encoder achieves the highest average accuracy (29.0%) among all single hard negative methods on MM-Math. Furthermore, while $\text{Retrieval}_{10}$ and $\text{Rule}_{10}$ negatives individually provide moderate gains, their combination further improves performance (29.4%), indicating their complementary nature. The best results of All Negatives are achieved when all three types of negatives are combined, demonstrating that diverse hard negatives collectively enhance the model’s geometric understanding. These findings underscore the importance of both the type and diversity of hard negatives in contrastive training for vision encoders.
Ablation Study 2: Image-based Negatives
Method #Num. GeoQA $M$-Math
Easy Med Hard Avg
Rule-Img 100K 52.9 44.0 25.4 1.0 28.3
Perturb-Img 4K 54.9 45.6 26.5 4.5 29.0
$\Delta$ (Perturb-Rule) -96K +2.0 +1.6 +1.1 +3.5 +0.7
Table 4: Effect of different image negative construction strategies. $\Delta$ rows give absolute gains of Perturb-Img over Rule-Img.
2.1. Code Perturbation vs. Rule-design: Assessing Image Negatives Construction Methods
We contrast two image–based negative construction methods. Perturb-Img applies our code perturbation method to 4K images, generating negative images that preserve similar structure while introducing subtle differences from the original diagrams. In contrast, Rule-Img is built from manually designed templates in the MAVIS corpus, enabling large-scale expansion. We construct 100K samples and retrieve most similar images as hard negatives. To ensure a fair comparison, we fix the positive-to-negative ratio at 1:10 in both settings. Results in Table 4 show that the 4K Perturb-Img setting exceeds the 100K Rule-Img on both GeoQA (54.9 vs. 52.9) and MM-Math (29.0 vs. 28.3). These results highlight that code-perturbed negatives, despite being smaller in scale, outperform synthetic data due to their semantic diversity and closer resemblance to real-world tasks.
2.2. Length/Angle Marked Matters? Cross-Modal Consistency Analysis
Figure 4: An example of vision encoder training data with or without numeric markings.
Visual Marking GeoQA $M$-MATH
Easy Med Hard Avg
w/ length/angle/etc. 54.9 45.6 26.5 4.5 29.0
w/o length/angle/etc. 51.7 50.0 25.1 4.5 28.6
Table 5: Impact of numeric markings on vision encoder training. w/ and w/o denote settings with and without such numeric annotations, respectively.
[…]
2.3. Hard Negative Performance Gains or Trade-off for In-Batch Training
Negative Strategy $M$-MATH
Easy Med Hard Avg
In-Batch 41.8 23.7 4.5 25.5
+image-based NS 51.7 (+9.9↑) 29.0 (+5.3↑) 4.5 32.1 (+6.6↑)
Table 6: Comparison of In-Batch and In-Batch training with the proposed image-based negative strategy. Mean score over 5 random seeds on 4K image-based negatives.
We leverage the proposed image-based negative strategy for the in-batch training to verify the effectiveness. In the in-batch negatives setting, each sample in the batch serves as a hard negative for others, with traditional in-batch typically selected at random. For a fair comparison, both methods use 10 negative examples per positive sample. As shown in Table 6, in-batch training with our proposed image-based negative strategy outperforms In-Batch on the easy, medium, and average subsets under the same amount of training data. This highlights the advantage of explicitly constructed image-based negatives in guiding the model to better distinguish visually similar examples, thus achieving higher overall accuracy. Furthermore, these results indicate that carefully selected mutual hard negatives within batches lead to more effective contrastive learning.
Ablation Study 3: Training Strategy and Robustness
Figure 5: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative ratio.
3.1. Data Scaling: More Hard Negatives Are Not Always Better
Our proposed MMGeoLM method allows flexible scaling of negatives. However, whether performance improves proportionally with an increasing number of hard negatives remains unclear. To address this, we investigate the impact of varying the number of retrieval-based negatives per sample from 5 to 50. As shown in Figure 5, MMGeoLM’s performance on MM-Math improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point. These results indicate that hard negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
3.2. Robustness to Geometric Image Transformations
Transformers Negatives $M$-MATH
Easy Med Hard Avg
No-Changing Text-based 48.6 29.1 9.0 31.9
[…]
Downscaling Text-based 45.6 26.8 0.0 28.4
No-Changing Image-based 45.6 26.5 4.5 29.0
[…]
Downscaling Image-based 48.2 25.6 4.5 28.7
Table 7: Robustness evaluation of MMGeoLM under image transformers (rotation and downscaling).
Diagrams in real exams frequently undergo transformations, such as rotation and scaling. To evaluate MMGeoLM’s robustness to these variations, we apply two image perturbations: (1) a $30^{\circ}$ clockwise rotation, and (2) downscaling images by a factor of 0.5. Table 7 summarizes the results under both text-based and image-based negative conditions. Model performance remains largely stable under a $30^{\circ}$ rotation, showing only a minor accuracy drop of 0.9 (31.9 $\rightarrow$ 31.0) in the text-based setting. This suggests the model has adequate robustness against moderate image rotations. In contrast, downscaling significantly impacts accuracy, dropping text-based performance from 31.9 to 28.4. This decline indicates that textual captions are particularly sensitive to reduced image resolution whereas the image-based negative demonstrates stronger robustness, benefiting from diagram scale variations.
Related Work
Mathematics-related research has recently received significant attention in LLMs and LMMs. For text-only mathematical reasoning, several works employ external tools like Tora (Gou et al. 2023) or intermediate step decomposition methods such as MAmmoTH (Yue et al. 2023), Metamath (Yu et al. 2023) and Math-Shepherd (Yue et al. 2023). Many multimodal mathematics benchmarks have been proposed, including mathematical competition-oriented OlympiadBench (He et al. 2024), geometry-focused datasets like Geometry3K (Lu et al. 2021), VisScience (Jiang et al. 2024), UniGeo (Chen et al. 2022a) and GeoQA (Chen et al. 2022a). Many works take effort on training multimodel math, including G-llava (Gao et al. 2023), Meta-LLaVA (Shi et al. 2024) and MAVIS (Zhang et al. 2024b). However, these models often struggle with fine-grained recognition of geometric elements.
Vision encoders are critical for capturing spatial and structural information in multimodal tasks. Prior works (Zhang, Awal, and Agrawal 2024; Doveh et al. 2023a, b; Singh et al. 2023) enhance image understanding using negative captions, but rarely target fine-grained geometric recognition. Methods such as NegCLIP (Yuksekgonul et al. 2022) and TriCLIP (Yang et al. 2024) focus on learning from negative samples, yet their effectiveness is limited by small or randomly selected negatives. In this work, we propose a scalable and geometry-aware negative learning strategy that explicitly constructs task-relevant hard negatives to improve geometric understanding.
Conclusion
This paper introduced two types of hard negatives, image-based and text-based negatives, targeted at geometric element understanding to enhance LMMs’ geometric reasoning. Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks. This work demonstrates the power of hard negative contrastive learning for improving LMMs with fine-grained visual understanding. Extensive ablation studies confirm the effectiveness of hard negatives, highlighting the crucial role advancing in multimodal mathematical reasoning.